Without using derivative find the maximum or minimum value of f(x) = |x+5| + 3

Grade:Upto college level

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = |x+5| + 3
f(x) = x+8, x \geq -5
f(x) = -x-2, x < -5
If you plot the graph of this function, you will see that the minimun value of f(x) is ‘3’ & maximum value is infinite.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty